A section of a stadium has 24 seats in the first row, 26 seats in the second row, increasing by 2 seats each row for a total of 33 rows. How many seats are in this section of the stadium.
-I can't seem to get this one right.

Answer by edjones(8007)

24+(32*2)=88 seats in the 33rd row.
(88+24)/2=56 seats average in each row.
56*33=1848 seats in this section of the stadium.
.
Ed